# Clust-PSI-PFL: A Population Stability Index Approach for Clustered Non-IID Personalized Federated Learning

### v1

@@ IV-A Experimental Design
Our experiments were conducted on a Linux workstation (Beastbianv2, kernel 6.1.0-20-amd64) equipped with an AMD Ryzen Threadripper PRO 5995WX processor (64 cores at 3.44 GHz), 500.0 GB of RAM, 222 GB of local storage, and an NVIDIA RTX A6000 GPU (48 GB). We used Python 3.10, and implemented FL baselines with the Flower [beutel2020flower] and FedLab [zeng2023fedlab] frameworks.
@@
We assess Clust-PSI-PFL across a broad suite of datasets spanning tabular, image, and text modalities, using diverse models and partition protocols. The specifications of the centralized datasets are summarized in Table I. These datasets were selected for their widespread use in prior studies on non-IID phenomena in FL. We employ two widely used partitioning protocols to simulate varying client data distributions. Dirichlet [jimenez2024fedartml] controls non-IID data with a single parameter $\alpha$: a smaller $\alpha$ generates more non-IID client partitions. In addition, we use the Similarity partition protocol [wang2023distribution], which is governed by a single parameter $S\!\in\![0,1]$: first, allocate $S*100\%$ of the dataset uniformly at random to clients (IID component); then sort the remaining $(100{-}S*100)\%$ by label and distribute it evenly across clients (label-skewed component). Larger $S$ produces more IID partitions, while smaller $S$ induces pathological non-IID data. TABLE I: Summary of datasets properties Dataset Modality Task # classes # examples ACSIncome [ding2021retiring] Tabular Classify high/low salary 2 1,664,500 Serengeti [swanson2015snapshot] Tabular Detect animal type 13 286,586 FMNIST [xiao2017fashion] Image Detect clothing type 10 70,000 CIFAR10 [krizhevsky2009learning] Image Classify objects 10 60,000 Sent140 [go2009twitter] Text Sentiment analysis 3 1,600,000 Amazon reviews [zhang2015character] Text Product classification 6 50,000
@@
For ACSIncome, we use a single linear layer (equivalent to logistic regression). For Serengeti, we employ a multilayer perceptron with three fully connected hidden layers (500 units each, ReLU activations) and a softmax output layer. For FMNIST, we adopt a convolutional neural network (CNN) with three convolutional layers (8, 16, and 32 channels), followed by max pooling and a fully connected layer with 2048 ReLU units. For CIFAR-10, we use a deeper CNN with three convolutional blocks; each block stacks two $3\times 3$ convolution layers (ReLU, same padding) with batch normalization, followed by $2\times 2$ max pooling and dropout (rates $0.2/0.3/0.4$ across the three blocks). The blocks use 32, 64, and 128 filters, respectively. The resulting features are flattened and passed to a 128-unit ReLU dense layer with batch normalization and dropout of 0.5, followed by a softmax layer over $C$ classes (input shape $H\times W\times 3$). TABLE II: Hyperparameters description and selection Baseline Hyperparameter Grid range Best Value All Learning rate {0.001,0.01,0.1} 0.001 All Batch size {16,32,64} 32 All Optimizer {Adam, SGD} Adam FedProx $\mu$ {0.001,0.01,0.1} 0.01 FedAvgM Momentum {0.7,0.9,0.99} 0.7 FedAdagrad $\tau$ {0.01,0.1} 0.1 FedAdagrad $\eta$ {0.1,0.3162} 0.3162 FedAdagrad $\eta_{l}$ {0.01,1} 1 FedYogi $\tau$ {0.001,0.01,0.1} 0.01 FedYogi $\eta$ {0.001,0.01} 0.001 FedYogi $\eta_{l}$ {0.01,1} 1 FedYogi $\beta_{1}$ {0.9,0.99} 0.9 FedYogi $\beta_{2}$ {0.9,0.99} 0.99 FedAdam $\tau$ {0.001,0.01,0.1} 0.001 FedAdam $\eta$ {0.001,0.01} 0.001 FedAdam $\eta_{l}$ {0.01,1} 1 FedAdam $\beta_{1}$ {0.9,0.99} 0.9 FedAdam $\beta_{2}$ {0.9,0.99} 0.99 PoC d {6,8,10,16} 8 HACCS $\rho$ {0.1,0.5,0.95} 0.95 FedCLS $\tau$ {0.1,0.5,0.95} 0.1 CFL $\epsilon_{1}$ {0.00001,0.0001,0.001} 0.00001 CFL $\epsilon_{2}$ {0.001, 0.01, 0.1} 0.1 CFL $\gamma$ {0.1,0.5,0.9} 0.5 FedSoft $n$ {5,10,50} 5 Clust-PSI-PFL (ours) $\tau$ {From 2 to K1K-1} Optimal $\tau$ For Sent140, we utilized a sequential NN with an embedding layer, a dropout layer (rate 0.5), an LSTM layer (10 units, dropout 0.2, recurrent dropout 0.2), and a dense output layer with softmax activation. For Amazon reviews, we used a lightweight deep neural network (DNN): a trainable embedding layer (vocabulary size $V$, embedding dimension $d$) followed by a Flatten layer, a small dense hidden layer with 8 ReLU units, and a softmax output over $C$ classes. We run the experiments using five different data partitions every time (i.e., five random seeds) to improve robustness. With standard deviations mostly below 0.02 (see Table III and IV), additional seeds add little value; five were enough to obtain stable estimates and preserve method rankings while avoiding unnecessary computation. The candidate hyperparameter grids and the corresponding optimal settings for each baseline are summarized in Table II. The search ranges were chosen in accordance with recommendations from the original baseline papers. Final values were selected via an exhaustive grid search over all parameter combinations. We train for $T=40$ communication rounds; at each round, a fraction $q=0.5$ of the $K$ clients is sampled uniformly without replacement, and each selected client runs $E=5$ local epochs. We use official test sets when available; otherwise, we adopt a train/test split of $80\%/20\%$.
@@
This section describes the set of metrics that are often considered in our experiments. Accuracy. The fraction of correctly classified samples relative to the total number of samples; larger values indicate better model performance. Because data in FL are partitioned across clients, accuracy can be evaluated at two levels: global (after aggregating across clients) and local (per client). Local accuracy. Formally, local accuracy is computed individually for each client $k$ as follows: $A_{k}=\frac{CC_{k}}{n_{k}}$ where $CC_{k}$ indicates the number of correctly classified samples on client $k$. Global accuracy. The global accuracy is calculated as the weighted average of local client accuracies, where the weight corresponds to the number of samples that each client holds: $A_{global}=\frac{\sum_{k=1}^{K}n_{k}A_{k}}{\sum_{k=1}^{K}n_{k}}$ with $K$ the total number of participating clients and $n_{k}$ the number of data samples on client $k$. Client fairness. We measure the difference of each solution from 1.0 (perfect accuracy). For such purpose, we calculate the average distance $AD=\frac{1}{K}\sum_{i=1}^{K}(|A_{i}-1.0|)$ and the corresponding standard deviation as $SDAD=\sqrt{\frac{1}{K}\sum_{i=1}^{K}(|A_{i}-1.0|-AD)^{2}}$. Smaller AD values represent a smaller distance to the perfect model (desired), and smaller SDAD showcases a smaller variation (desired) in such distances.
@@ V Limitations of Clust-PSI-PFL
While Clust-PSI-PFL shows consistent gains across datasets and non-IID scenarios, our current implementation does not incorporate privacy-preserving mechanisms such as DP or MPC. Integrating MPC-based secure aggregation does not change model performance (it adds only runtime/communication overhead), whereas DP may affect performance, with the impact governed by the budget $\epsilon$. Second, the present version of Clust-PSI-PFL focuses on label skew: PSI features are derived from class frequency, and clients are clustered by similarity in label proportions. Other non-IID data types (e.g., attribute skew, quantity imbalance) are not explicitly modeled; thus, when labels are balanced but attribute distributions or data volumes differ, clusters may be suboptimal. We view this as a scoped design choice rather than a fundamental limitation: the PSI machinery can be extended to discretized attributes, marginals, or joint label–attribute bins, and incorporate sample counts.
@@ VI Conclusion and Future Work
In this study, we present Clust-PSI-PFL, a PFL framework that leverages the PSI as a principled metric for client clustering under non-IID data conditions. We demonstrate that PSI reliably quantifies distributional non-IID data across tabular, image, and text modalities, enabling the formation of distributionally coherent client groups and, in turn, yielding substantial gains in both global and local performance, even in highly non-IID regimes. Relative to state-of-the-art baselines, Clust-PSI-PFL achieves up to 18% higher global accuracy while enhancing client fairness, resulting in a relative improvement of 37%. These results position Clust-PSI-PFL as a practical mechanism for robust decentralized learning in non-IID data federated environments. In future work, we will examine the applicability of PSI beyond label skew to other forms of data skew, such as attribute skew and quantity skew. Moreover, integrating formal privacy mechanisms, such as DP[erlingsson2014rappor] and MPC[bohler2021secure], to strengthen the confidentiality of our FL pipeline lies outside the scope of this study and remains a promising direction for follow-up research.
VII Acknowledgements
Daniel M. Jimenez-Gutierrez and Andrea Vitaletti were partially supported by PNRR351 TECHNOPOLE – NEXT GEN EU Roma Technopole – Digital Transition, FP2 – Energy transition and digital transition in urban regeneration and construction and Sapienza Ateneo Research grant “La disintermediazione della Pubblica Amministrazione: il ruolo della tecnologia blockchain e le sue implicazioni nei processi e nei ruoli della PA.” Aris Anagnostopoulos was supported by the ERC Advanced Grant 788893 AMDROMA, the EC H2020RIA project “SoBigData++” (871042), the PNRR MUR project PE0000013-FAIR, the PNRR MUR project IR0000013-SoBigData.it, and the MUR PRIN project 2022EKNE5K “Learning in Markets and Society.” Ioannis Chatzigiannakis was supported by PE07-SERICS (Security and Rights in the Cyberspace) – European Union Next-Generation-EU-PE0000014 (Piano Nazionale di Ripresa e Resilienza – PNRR). Andrea Vitaletti was supported by PE11 - MICS (Made in Italy – Circular and Sustainable) – European Union Next-Generation-EU (Piano Nazionale di Ripresa e Resilienza – PNRR) and the project SERICS (PE00000014) under the MUR National Recovery and Resilience Plan funded by the European Union - NextGenerationEU.

### v2

@@ IV-A Experimental Design
All experiments ran on a single Linux machine (Beastbianv2, kernel 6.1.0-20-amd64) with an AMD Threadripper PRO 5995WX CPU (64 cores), 500 GB RAM, and an NVIDIA RTX A6000 GPU (48 GB). The codebase was implemented in Python 3.10 and relied on Flower [3] and FedLab [37] for the FL baselines.
@@
We assess Clust-PSI-PFL across a broad suite of datasets spanning tabular, image, and text modalities, using diverse models and partition protocols. The specifications of the centralized datasets are summarized in Table I. These datasets were selected for their widespread use in prior studies on non-IID phenomena in FL. We employ two widely used partitioning protocols to simulate varying client data distributions. Dirichlet [15] controls non-IID data with a single parameter $\alpha$: a smaller $\alpha$ generates more non-IID client partitions. In addition, we use the Similarity partition protocol [32], which is governed by a single parameter $S\!\in\![0,1]$: first, allocate $S*100\%$ of the dataset uniformly at random to clients (IID component); then sort the remaining $(100{-}S*100)\%$ by label and distribute it evenly across clients (label-skewed component). Larger $S$ produces more IID partitions, while smaller $S$ induces pathological non-IID data. TABLE I: Summary of datasets properties Dataset Modality Task # classes # examples ACSIncome [7] Tabular Classify high/low salary 2 1,664,500 Serengeti [29] Tabular Detect animal type 13 286,586 FMNIST [34] Image Detect clothing type 10 70,000 CIFAR10 [18] Image Classify objects 10 60,000 Sent140 [10] Text Sentiment analysis 3 1,600,000 Amazon reviews [38] Text Product classification 6 50,000
@@
For ACSIncome, we train a shallow classifier consisting of a single fully connected layer with a softmax/sigmoid output (i.e., a logistic-regression baseline). For Serengeti, we use a feed-forward network with three hidden layers of 500 neurons each and ReLU non-linearities, ending with a softmax classification head. For FMNIST, we rely on a lightweight CNN made of three convolutional stages with 8, 16, and 32 feature maps, respectively; the convolutional stack is followed by max-pooling and a dense layer with 2048 ReLU units before the final classifier. For CIFAR-10, we adopt a deeper convolutional architecture composed of three feature-extraction blocks. In each block, two $3\times 3$ convolutions (ReLU, same padding) are applied and regularized with batch normalization, after which a $2\times 2$ max-pooling operation reduces spatial resolution. We further apply dropout after each block, increasing the rate from 0.2 to 0.3 and 0.4 in the successive blocks, while using 32, 64, and 128 filters in the first, second, and third blocks, respectively. The final activation maps are flattened and fed to a 128-unit fully connected ReLU layer with batch normalization and dropout (p=0.5p{=}0.5), and a softmax output over the $C$ classes (input size $H\times W\times 3$). TABLE II: Hyperparameters description and selection Baseline Hyperparameter Grid range Best Value All Learning rate {0.001,0.01,0.1} 0.001 All Batch size {16,32,64} 32 All Optimizer {Adam, SGD} Adam FedProx $\mu$ {0.001,0.01,0.1} 0.01 FedAvgM Momentum {0.7,0.9,0.99} 0.7 FedAdagrad $\tau$ {0.01,0.1} 0.1 FedAdagrad $\eta$ {0.1,0.3162} 0.3162 FedAdagrad $\eta_{l}$ {0.01,1} 1 FedYogi $\tau$ {0.001,0.01,0.1} 0.01 FedYogi $\eta$ {0.001,0.01} 0.001 FedYogi $\eta_{l}$ {0.01,1} 1 FedYogi $\beta_{1}$ {0.9,0.99} 0.9 FedYogi $\beta_{2}$ {0.9,0.99} 0.99 FedAdam $\tau$ {0.001,0.01,0.1} 0.001 FedAdam $\eta$ {0.001,0.01} 0.001 FedAdam $\eta_{l}$ {0.01,1} 1 FedAdam $\beta_{1}$ {0.9,0.99} 0.9 FedAdam $\beta_{2}$ {0.9,0.99} 0.99 PoC d {6,8,10,16} 8 HACCS $\rho$ {0.1,0.5,0.95} 0.95 FedCLS $\tau$ {0.1,0.5,0.95} 0.1 CFL $\epsilon_{1}$ {0.00001,0.0001,0.001} 0.00001 CFL $\epsilon_{2}$ {0.001, 0.01, 0.1} 0.1 CFL $\gamma$ {0.1,0.5,0.9} 0.5 FedSoft $n$ {5,10,50} 5 Clust-PSI-PFL (ours) $\tau$ {From 2 to K1K-1} Optimal $\tau$ For Sent140, we employ a compact sequence model consisting of an embedding layer, dropout (p=0.5p{=}0.5), a 10-unit LSTM (dropout/recurrent dropout $0.2$), and a softmax classifier. For Amazon Reviews, we use a small embedding-based DNN (embedding $\rightarrow$ flatten $\rightarrow$ 8-unit ReLU layer $\rightarrow$ softmax). All results are averaged over five independent partitions (five seeds). With standard deviations mostly below 0.02 (see Table III and IV), additional seeds add little value; five were enough to obtain stable estimates and preserve method rankings while avoiding unnecessary computation. The candidate hyperparameter grids and the corresponding optimal settings for each baseline are summarized in Table II. The search ranges were chosen in accordance with recommendations from the original baseline papers. Final values were selected via an exhaustive grid search over all parameter combinations. We train for $T=40$ communication rounds; at each round, a fraction $q=0.5$ of the $K$ clients is sampled uniformly without replacement, and each selected client runs $E=5$ local epochs. We use official test sets when available; otherwise, we adopt a random train/test split of $80\%/20\%$.
@@
This section summarizes the evaluation metrics used in our experiments. Accuracy. We report the proportion of correctly classified samples; higher values correspond to better predictive performance. Since FL data are distributed, we consider both per-client and federation-level accuracy. Local accuracy. For each client $k$, we compute: $A_{k}=\frac{CC_{k}}{n_{k}}$ where $CC_{k}$ is the number of correct predictions on client $k$ and $n_{k}$ is its sample count. Global accuracy. We obtain a single federation score by averaging local accuracies weighted by client data sizes: $A_{global}=\frac{\sum_{k=1}^{K}n_{k}A_{k}}{\sum_{k=1}^{K}n_{k}}$ with $K$ the total number of participating clients and $n_{k}$ the number of data samples on client $k$. Client fairness. To characterize performance disparities across clients, we measure each client’s deviation from perfect accuracy (1.0) and summarize it via the mean distance $AD=\frac{1}{K}\sum_{i=1}^{K}(|A_{i}-1.0|)$ and its dispersion $SDAD=\sqrt{\frac{1}{K}\sum_{i=1}^{K}(|A_{i}-1.0|-AD)^{2}}$. Lower $AD$ and $SDAD$ indicate, respectively, higher overall fairness and smaller variability in client outcomes.
@@ V Limitations of Clust-PSI-PFL
While Clust-PSI-PFL shows consistent gains across datasets and non-IID scenarios, our current implementation does not incorporate privacy-preserving mechanisms such as DP or MPC. Integrating MPC-based secure aggregation does not change model performance (it adds only runtime/communication overhead), whereas DP may affect performance, with the impact governed by the budget $\epsilon$. Providing formal privacy guarantees is left to DP/MPC-enhanced variants. Second, the present version of Clust-PSI-PFL focuses on label skew: PSI features are derived from class frequency, and clients are clustered by similarity in label proportions. Other non-IID data types (e.g., attribute skew, quantity imbalance) are not explicitly modeled; thus, when labels are balanced, but attribute distributions or data volumes differ, clusters may be suboptimal. We view this as a scoped design choice rather than a fundamental limitation: the PSI machinery can be extended to discretized attributes, marginals, or joint label–attribute bins, and incorporate sample counts. Finally, when some clients have extremely small local datasets, their label histograms (and thus PSI) can be noisy. A simple mitigation is to down-weight such clients during clustering or append $n_{i}$ as an auxiliary clustering feature. A systematic evaluation of such extremely small-client regimes is an important direction for future work.
@@ VI Conclusion and Future Work
In this study, we present Clust-PSI-PFL, a PFL framework that leverages the PSI as a principled metric for client clustering under non-IID data conditions. We demonstrate that PSI reliably quantifies distributional non-IID data across tabular, image, and text modalities, enabling the formation of distributionally coherent client groups and, in turn, yielding substantial gains in both global and local performance, even in highly non-IID regimes. Relative to state-of-the-art baselines, Clust-PSI-PFL achieves up to 18% higher global accuracy while enhancing client fairness, resulting in a relative improvement of 37%. These results position Clust-PSI-PFL as a practical mechanism for robust decentralized learning in non-IID data federated environments. In future work, we will examine the applicability of PSI beyond label skew to other forms of data skew, such as attribute skew and quantity skew. Moreover, integrating formal privacy mechanisms, such as DP[9] and MPC[4], to strengthen the confidentiality of our FL pipeline lies outside the scope of this study and remains a promising direction for follow-up research.
Appendix: Artifact Description/Artifact Evaluation Artifact Description (AD)
VII Overview of Contributions and Artifacts
VII-A Paper’s Main Contributions
$C_{1}$ PSI-based quantification of non-IIDness in FL and comparison to alternative metrics. $C_{2}$ Clust-PSI-PFL: PSI-guided client clustering (K-means++) with silhouette-based cluster count selection and cluster-wise personalization. $C_{3}$ Empirical validation across 6 datasets and 2 partition protocols (Dirichlet($\alpha$), Similarity($S$)), reporting accuracy and fairness improvements vs. baselines.
VII-B Computational Artifacts
$A_{1}$ Code repository (implementation + experiments):
https://github.com/Sapienza-University-Rome/clust_psi_pfl
TABLE V: Artifact-to-paper mapping (compact). ID Supports Reproduces $A_{1}$ $C_{1}$$C_{3}$ Main experimental results (accuracy and fairness) and clustering behavior reported in the paper; see README.md for runnable configurations.
VIII Artifact Identification (A1)
VIII-A Computational Artifact $A_{1}$
Relation To Contributions
$A_{1}$ is the reference implementation of Clust-PSI-PFL and the baselines used in the evaluation (methods selectable via CLI). It enables (i) PSI-based non-IID quantification, (ii) PSI-driven clustering and $\tau$ selection, and (iii) federated training/evaluation on the supported datasets and partition protocols. License: MIT.
Expected Results
Expected outcomes (trend-level): (i) Stronger non-IID (smaller $\alpha$ / smaller $S$) yields worse baseline performance; (ii) Clust-PSI-PFL improves global accuracy and reduces client disparity (fairness); (iii) Clustering selects a small $\tau$ under the silhouette criterion. Exact numbers may vary due to nondeterminism and environmental differences.
Expected Reproduction Time (in Minutes)
Setup: 10–30 min. Execution: 30–60 min (smoke test); paper-like sweeps: 26 hours, including all the baselines for the slower dataset, which is Sent140. Analysis: 10–20 min.
Artifact Setup (incl. Inputs)
Hardware
Validated on Linux with high-end CPU/RAM and an NVIDIA RTX A6000 (48GB). A CUDA GPU is recommended for image/text datasets; CPU-only reproduction is possible but slower. Disk: 1TB.
Software
Linux; Python 3.10 via conda env clust_psi_pfl_py310. Install all dependencies from environment.yml. Key versions: Follow the versions included in the environment.yml file.
Datasets / Inputs
Supported datasets: acs_income, serengeti, fmnist, cifar10, sent140, amazon_reviews. Sent140 must be downloaded manually (not shipped); follow the README.md instructions and expected folder layout under data/sent140/. For the remaining datasets, no manual dataset folder preparation is required; in particular, cifar10 and fmnist are obtained via standard dataset utilities (i.e., no repository data folder is needed in advance).
Installation and Deployment
Clone repo, then create/activate conda env from environment.yml. If running Sent140, download and place required files under data/sent140/ (README.md). It can be run on a single machine (GPU is optional but recommended, especially for image and text datasets).
Artifact Execution
Workflow (tasks/dependencies): $T_{1}$ partition data (--dataset, --partitioner, --non-iid-param, --num-clients) $\rightarrow$ $T_{2}$ (Clust-PSI-PFL only) compute PSI features + cluster selection (silhouette) / fixed --tau-clusters $\rightarrow$ $T_{3}$ federated training/evaluation (--agg-method) producing accuracy and fairness metrics. For non-clustering baselines: $T_{1}\rightarrow T_{3}$. Paper-like settings: T=40T\!=\!40 rounds, E=5E\!=\!5 local epochs, 5 seeds/partitions; client fraction q=0.5q\!=\!0.5.
Artifact Analysis (incl. Outputs)
Outputs include global accuracy, per-client local accuracy, and fairness metrics ($AD$, SDADSDAD), printed.
Artifact Evaluation (AE)
VIII-A Computational Artifact $A_{1}$
Artifact Setup (incl. Inputs)
https://github.com/Sapienza-University-Rome/clust_psi_pfl/blob/main/README.md Install/Run (minimal). Create env: conda env create -f environment.yml; activate: conda activate clust_psi_pfl_py310. Smoke test: python main.py. If using Sent140: download into data/sent140/ as per README.md.
Artifact Execution
[…]
Artifact Analysis (incl. Outputs)
Expected results: Under strong non-IID (small $\alpha$ / $S$), Clust-PSI-PFL should improve global accuracy and reduce client disparity vs. baselines (trend-level agreement). Evaluation: Compare reproduced accuracy/fairness trends with the corresponding results in the paper (same dataset/partition). Acceptable tolerance: $\pm 1$ percentage point for global accuracy and $\pm 5\%$ (relative) for fairness metrics ($AD$, SDADSDAD); trend-level agreement across non-IID settings ($\alpha$/$S$) is required. Contribution link: PSI/non-IID trends support $C_{1}$; clustering behavior supports $C_{2}$; accuracy+fairness improvements support $C_{3}$.